# Visual CoT: Advancing Multi-Modal Language Models with a Comprehensive Dataset and Benchmark for Chain-of-Thought Reasoning

### v2

@@ Abstract
Multi-Modal Large Language Models (MLLMs) have demonstrated impressive performance in various VQA tasks. However, they often lack interpretability and struggle with complex visual inputs, especially when the resolution of the input image is high or when the interested region that could provide key information for answering the question is small. To address these challenges, we collect and introduce the large-scale Visual CoT dataset comprising 438k question-answer pairs, annotated with intermediate bounding boxes highlighting key regions essential for answering the questions. Additionally, about 98k pairs of them are annotated with detailed reasoning steps. Importantly, we propose a multi-turn processing pipeline that dynamically focuses on visual inputs and provides interpretable thoughts. We also introduce the related benchmark to evaluate the MLLMs in scenarios requiring specific local region identification. Extensive experiments demonstrate the effectiveness of our framework and shed light on better inference strategies. The Visual CoT dataset, benchmark, and pre-trained models are released to foster further research in this direction.
@@ 1 Introduction
However, many popular MLLMs are primarily trained to respond to instructions based on visual inputs, employing a decoder-only autoregressive design as a single black box. While these models exhibit impressive generation capabilities, they suffer from inaccurate information [30] and even hallucinations [15]. Moreover, the black-box design hinders the interpretability of visual-language models. Additionally, the potential of multi-turn in-context capability and the advantages of chain-of-thought [62, 75, 69] for LLMs have not been extensively explored in MLLMs. Some recent works, such as multimodal-CoT [76] and [68, 67], have shown improvements by incorporating text-level chain-of-thought reasoning or in-context learning. However, it remains uncharted whether existing MLLMs can benefit from chain-of-thought reasoning in the visual understanding process, along with their interpretability remains largely unexplored.
@@ 3.1 Data Generation
Text/Doc. We choose five text-related datasets to create data in this domain: TextVQA [53], DocVQA [43], DUDE [57], TextCaps [52], SROIE [17]. The five datasets focus on text recognition and comprehension in a variety of images and documents. TextVQA, DocVQA, DUDE and SROIE have already provided question-answer pairs, which we directly adopt. TextCaps, providing only captions and OCR tokens, required us to employ a linguistic annotator to create corresponding questions and answers (see further details in Appendix E.1.) For the visual CoT bboxes, we then apply PaddleOCR[12] to detect OCR-identified regions in the image, and specify the CoT bounding boxes as the region that consists of words and sentences aligning with the answer. Furthermore, we also design a filtering pipeline to improve content quality. This process ensures that the areas highlighted by the bounding boxes are directly relevant to the questions.
@@ D.1 Performance evaluation
In Tab. 8 and Tab. 9, we showcase the baseline performance of our model, where it directly answers questions without employing the visual CoT process. Multi-modal Large Language Models Benchmarks. In Tab. 8, we evaluate our model on recently proposed MLLM benchmarks such as MME [13], POPE [30], MMbench [36], ScienceQA [37], TextVQA [53], GQA [18]. Our model still achieves comparative results across all benchmarks. This performance indicates that the visual CoT data we proposed not only enhances visual comprehension in CoT-specific scenarios but also boosts the model’s overall visual understanding in standard inference setups. As demonstrated in Tab. 10, the implementation of visual CoT enables our model to achieve superior performance even with a lower resolution and a reduced number of visual tokens. This finding highlights the efficiency and effectiveness of the visual CoT approach in enhancing model accuracy.
Visual grounding. Furthermore, we evaluate VisCoT on REC benchmarks with RefCOCO [19], RefCOCO+ [41], and RefCOCOg [41] datasets. Our model outperforms the previous state-of-the-art models, including the specialist models such as G-DINO-L [35] and UNINEXT [66]. Notably, even with a minimal setup (7B LLM & 224 resolution), our approach outperforms methods that utilize higher resolutions or larger LLM models. This demonstrates that our dataset, enhanced with intermediate bounding boxes, significantly improves the model’s precision in locating and understanding referred objects or regions.

### v3

@@ Abstract
Multi-Modal Large Language Models (MLLMs) have demonstrated impressive performance in various VQA tasks. However, they often lack interpretability and struggle with complex visual inputs, especially when the resolution of the input image is high or when the interested region that could provide key information for answering the question is small. To address these challenges, we collect and introduce the large-scale Visual CoT dataset comprising 438k question-answer pairs, annotated with intermediate bounding boxes highlighting key regions essential for answering the questions. Additionally, about 98k pairs of them are annotated with detailed reasoning steps. Importantly, we propose a multi-turn processing pipeline that dynamically focuses on visual inputs and provides interpretable thoughts. We also introduce the related benchmark to evaluate the MLLMs in scenarios requiring specific local region identification. Extensive experiments demonstrate the effectiveness of our framework and shed light on better inference strategies. The Visual CoT dataset, benchmark, and pre-trained models are available on this webpage to support further research in this area.
00footnotetext: Corresponding author.
@@ 1 Introduction
However, many popular MLLMs [47, 58, 23, 85, 7, 9, 76, 75, 83] and related benchmarks [35, 8, 22, 73, 74] are primarily trained to respond to instructions based on visual inputs, employing a decoder-only autoregressive design as a single black box. While these models exhibit impressive generation capabilities, they suffer from inaccurate information [36] and even hallucinations [18]. Moreover, the black-box design hinders the interpretability of visual-language models. Additionally, the potential of multi-turn in-context capability and the advantages of chain-of-thought [70, 89, 81] for LLMs have not been extensively explored in MLLMs. Some recent works, such as multimodal-CoT [90] and [80, 79], have shown improvements by incorporating text-level chain-of-thought reasoning or in-context learning. However, it remains uncharted whether existing MLLMs can benefit from chain-of-thought reasoning in the visual understanding process, along with their interpretability remains largely unexplored.
@@ 3.1 Data Generation
Text/Doc. We choose five text-related datasets to create data in this domain: TextVQA [61], DocVQA [50], DUDE [65], TextCaps [60], SROIE [20]. The five datasets focus on text recognition and comprehension in a variety of images and documents. TextVQA, DocVQA, DUDE and SROIE have already provided question-answer pairs, which we directly adopt. TextCaps, providing only captions and OCR tokens, required us to employ a linguistic annotator to create corresponding questions and answers (see further details in Appendix E.1). For the visual CoT bboxes, we then apply PaddleOCR[15] to detect OCR-identified regions in the image, and specify the CoT bounding boxes as the region that consists of words and sentences aligning with the answer. Furthermore, we also design a filtering pipeline to improve content quality. This process ensures that the areas highlighted by the bounding boxes are directly relevant to the questions.
@@ 6 Conclusion
Acknowledgement. This project is funded in part by National Key R&D Program of China Project 2022ZD0161100, by the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK, by General Research Fund of Hong Kong RGC Project 14204021. Hongsheng Li is a PI of CPII under the InnoHK.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix F.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix G.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We have provided the related details in Appendix. The code, training data, benchmark, and checkpoints can be found in this GitHub repo: https://github.com/deepcs233/Visual-CoT
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See ‘Training Details’ in Section Experiments. We also provide reproducible scripts that contain all hyperparameters in this GitHub repo: https://github.com/deepcs233/Visual-CoT
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix B.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We provide our code, training data, and checkpoints in this GitHub repo: https://github.com/deepcs233/Visual-CoT
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix H.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix H.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ D.1 Performance evaluation
In Tab. 8 and Tab. 9, we showcase the baseline performance of our model, where it directly answers questions without employing the visual CoT process.
Multi-modal Large Language Models Benchmarks. In Tab. 8, we evaluate our model on recently proposed MLLM benchmarks such as MME [16], POPE [36], MMbench [42], ScienceQA [43], TextVQA [61], GQA [21]. Our model still achieves comparative results across all benchmarks. This performance indicates that the visual CoT data we proposed not only enhances visual comprehension in CoT-specific scenarios but also boosts the model’s overall visual understanding in standard inference setups. As demonstrated in Tab. 10, the implementation of visual CoT enables our model to achieve superior performance even with a lower resolution and a reduced number of visual tokens. This finding highlights the efficiency and effectiveness of the visual CoT approach in enhancing model accuracy.
Visual grounding. Furthermore, we evaluate VisCoT on REC benchmarks with RefCOCO [24], RefCOCO+ [48], and RefCOCOg [48] datasets. Our model outperforms the previous state-of-the-art models, including the specialist models such as G-DINO-L [41] and UNINEXT [78]. Notably, even with a minimal setup (7B LLM & 224 resolution), our approach outperforms methods that utilize higher resolutions or larger LLM models. This demonstrates that our dataset, enhanced with intermediate bounding boxes, significantly improves the model’s precision in locating and understanding referred objects or regions. “Top-1 Accuracy@0.5” refers to the accuracy of a model in predicting the correct bounding box as the top prediction when the Intersection over Union (IoU) between the predicted and ground truth bounding boxes meets or exceeds 50%.